# Fluidic FlowBots: Intelligence embodied in the characteristics of recirculating fluid flow

### v1

@@ Abstract
The one-to-one mapping of control inputs to actuator outputs results in elaborate routing architectures that limit how complex fluidic soft robot behaviours can currently become. Embodied intelligence can be used as a tool to counteract this phenomenon. Control functionality can be embedded directly into actuators by leveraging the characteristics of fluid flow phenomena. Whilst prior soft robotics work has focused exclusively on actuators operating in a state of transient/no flow (constant pressure), or pulsatile/alternating flow, our work begins to explore the possibilities granted by operating in the closed-loop flow recirculation regime. Here we introduce the concept of FlowBots: soft robots that utilise the characteristics of continuous fluid flow to enable the embodiment of complex control functionality directly into the structure of the robot. FlowBots have robust, integrated, no-moving-part control systems, and these architectures enable: monolithic additive manufacturing methods, rapid prototyping, greater sustainability, and an expansive range of applications. Based on three FlowBot examples: a bidirectional actuator, a gripper, and a quadruped swimmer - we demonstrate how the characteristics of flow recirculation contribute to simplifications in fluidic analogue control architectures. We conclude by outlining our design and rapid prototyping methodology to empower others in the field to explore this new, emerging design field, and design their own FlowBots.
@@ I Introduction
Whilst there as several other notable examples of previous fluidic soft robots that used “pulsatile”, or “cyclic” modes of flow (such as [35, 8, 36, 37, 19, 38]), a condition that has not yet been explored and well-characterised in the context of soft fluidic robots is a state of continuous flow recirculation. Figure 1 illustrates the difference between soft fluidic robots based on the “steady state”, “pulsatile/cyclic”, and “recirculating” modes of flow, as defined in this paper. Fluid can be recirculated in a system for example to achieve thermal regulation, as in the work of Cacucciolo et al. [39], and others [40, 41]. Liao et al. [41] showed that the pressure of this recirculating fluid can also be used to inflate the cavities of an actuator, achieving deformation whilst simultaneously taking advantage of the thermal regulation characteristic (which they used to modify the stiffness of the actuator).
@@
Figure 2: Inspiration for exploiting fluid flow phenomena as a form of embodied intelligence can be found in nature [43]. The geometry of the shark intestine (c) resembles that of a Tesla-valve (a, b). Flow circulation induced by the bellowed geometry results in a diode-like behaviour, making it harder for fluid to flow in one direction (left-to-right as pictured on the figure) than the other (right-to-left). This type of fluidic diode has no moving components; the added functionality is embodied into the structure of the device.
@@
This paper focuses on embedding additional functionality directly into the bodies of soft robots, by taking advantage of the properties of recirculating fluid flow. Here, we demonstrate that new types of soft robots can be manufactured as a single part, with embedded control, and actuation. This mode of manufacture enables rapid prototyping, increases robustness, reduces waste, and minimises the risk of environmental pollution (opening up opportunities for new applications in e.g. wildlife biodiversity monitoring). We term such robots “FlowBots”.
@@
The focus of this paper is to demonstrate the principles of operation of FlowBots based on three simple examples: a bidirectional actuator, a gripper, and a quadruped swimmer robot. Each example builds upon the operating principles of the previous one. The bidirectional actuator demonstrates how pressure asymmetries arising due to viscous losses can be used for analogue control using localised flow channel constrictions. The gripper shows how flow recirculation can simplify control architectures of multi-actuator systems due to the ability of switching between connection in parallel, and in series. The swimmer FlowBot builds upon the control architecture used for the gripper, and serves as a demonstration of how the embodiment of intelligence in fluid flow characteristics allows to create robots capable of complex behaviours, yet manufactured as single, robust components, without any moving parts. Section II describes some of our experimental work in evaluating the performance of FlowBots with air and water as the operating fluid. We conclude with a discussion on some of the design challenges associated with building FlowBots, empowering others in the field to build their own.
@@ II Results
If the location of the constriction was moved further upstream, or downstream of the tip, the deformation characteristic of the actuator would be permanently skewed to one side. Having the actuator deform differently in both directions at the same pressure difference could be desirable in certain applications. In this paper, all the actuators are symmetric, but we utilise external variable constrictions to regulate the direction of flow, and the pressure difference across the actuator. These constrictions can be implemented using additional valves, by manually blocking/kinking the outlet vents/tubes, or by using soft buttons (see the supplementary video). Figure 3 shows how we were able to use two such constrictions to achieve analogue control of the actuator in two directions, by regulating the pressure across it over a continuous spectrum. This analogue mapping can also be achieved by adjusting the direction of flow, as well as the flow rate of the supply pump (hydraulic case), or compressed air supply pressure (pneumatic case).
Figure 3: The principle of operation of the bidirectional actuator. By varying the outlet vent constrictions, the direction, and magnitude of the flow, as well as the pressure difference across the actuator can be controlled with an analogue characteristic. Reversing the flow direction causes the actuator to deform symmetrically in the other direction due to the mirroring of the pressure asymmetry. The percentage value of the constriction is an approximation; we estimated it indirectly by measuring the pressure difference across the actuator using two pressure gauge probes.
@@
Section II-A demonstrated how to control a single bidirectional actuator, two controls (vent outlet constrictions) are used. To fully control two such actuators, four controls would normally be needed. By exploiting the characteristics of recirculating fluid flow, we were able to reduce this fan-in to three. The simplification comes from the fact that under the state of recirculating flow, the actuators can operate not only in a parallel, but also a series connection. The ability to switch between the two configurations allows to simplify the control architecture, and opens up the design space for constructing FlowBots consisting of several actuators, that are capable of more complex behaviours.
@@
The control characteristics of FlowBots are embodied in the interaction of the fluid flow, and the structure of the robot structure. No added parts, or moving components are required, and, as a result, entire robots can be additively manufactured as a single part, in one print job. Figure 5 depicts the operation of a quadruped swimming robot that we have built based on the control architecture used for the gripper design described in Section II-B. As in the case of the gripper, we can fully control each pair of limbs for this FlowBot using only three control inputs rather than four, thanks to the ability to switch between parallel and series connection of the actuators. Despite the added complexity of the design, the entire 3D stack of the control architecture comes packaged into the body of the robot, and has been additively manufactured as a single component. After the print job is over, the robot is ready to use. Embodying intelligence in this way makes it easy to prototype FlowBots with ever more complex 3D designs.
Figure 5: Operation of a quadruped swimmer FlowBot that builds upon the control architecture used for the gripper described by Section II-B. Due to the embodiment of control into the body of the robot, as well as the characteristics of recirculating flow, FlowBots of increasing complexity can still be manufactured as single components.
@@
To account for the difference in the viscosity of the fluids, we used the pressure difference between the inlets of the actuator as the independent variable in the experiment. This approach allowed us to validate whether the actuator deforms the same amount with the same pressure difference, independent of whether air, or water, was used as the working fluid. We examined a range of pressures from 1.25 bar to 2.5 bar (in intervals of 0.25 bar), testing in both flow directions with 3 measurement repetitions per data point. We also repeated the measurements with another identical actuator specimen. Due to the symmetry resulting from placing the channel constriction at the tip of the actuator, we expect it to deform the same amount in both directions, as explained in Section II-A. For more details about the experimental setup, please refer to the supplementary material.
@@ III Discussion
The channels in the actuators featured in the experiment are as narrow as 0.6 mm, and variations on the order of hundreds of microns at different locations along the flow path can disturb the symmetry between the pressure distributions in both directions; such deviations can contribute to marginal differences in the deformation in two directions. In our case, these errors did not significantly affect the robot’s desired functionality, but it might be a factor to consider for those designing FlowBots with very specific precision requirements. In these instances, we recommend carrying out adequate testing to ensure that the desired level of precision has been achieved.
@@
This paper demonstrates how designers can come up with entirely new fluidic soft robot concepts by taking advantage of the characteristics of fluid flow. Constant recirculation of the working fluid comes with the cost of energy expenditure, but it allows to simplify control architectures (as demonstrated in Section II-B), and to rapidly prototype and additively manufacture entirely new FlowBot designs. For additional design considerations to take into account when adapting systems to use recirculating flow, please refer to the supplementary material, where we have discussed these points in further detail.
@@
The exact behaviour of the fluid inside of the internal channels is difficult to characterise analytically due to the turbulence arising as a result of the combination of bellow geometries, multi-directional pipe bends, and varying channel cross-sections. Whilst we do invite others specialising in the domain to help better characterise these phenomena, we encourage those interested in coming up with their own FlowBot designs to take advantage of Finite Element Method (FEM) software. In our design methodology, we have been able to utilise Computational Fluid Dynamics (CFD) to obtain an approximation of the internal flow field that was more than adequate enough to facilitate effective design exploration and validation. By combining CFD with structural simulations using Finite Element Analysis (FEA) software we were able to simultaneously explore how modifying parameters in the design affects the fluid flow in the control system, as well as the mechanical deformation of the actuators.
@@
With our chosen method of Fused Deposition Modelling (FDM) 3D printing, we were able to manufacture all of the FlowBots featured in this paper as one part, without any supports or additional post-processing needed. Once taken off the print bed, each FlowBot is ready to use, as it integrates actuation, control, and even fluidic fittings in its design. In spite of the complex bellow geometry of its actuators, and the multi-layered internal 3D control architecture, the swimmer FlowBot described in Section II-C took under 7 hours to print on a cheap, \$200 Artillery Sidewinder X1 printer using Recreus Filaflex 70A filament (as well as 82A for the specimens in Section II-D). This design approach is more environmentally sustainable, adds robustness and saves mass and space due to the reduced need for complex routing/manifolds that come with a pressurising/valving system.
@@ IV Conclusion
This paper introduced the concept of FlowBots, a new generation of soft robots that operate under a state of fluid flow recirculation. FlowBots are controlled by taking advantage of fluid flow characteristics, allowing for embodied intelligence. This work demonstrates the operation of three simple FlowBot systems: a bidirectional actuator, a gripper, and a quadruped swimmer robot.

### v2

@@ Abstract
The one-to-one mapping of control inputs to actuator outputs results in elaborate routing architectures that limit how complex fluidic soft robot behaviours can currently become. Embodied intelligence can be used as a tool to counteract this phenomenon. Control functionality can be embedded directly into actuators by leveraging the characteristics of fluid flow phenomena. Whilst prior soft robotics work has focused exclusively on actuators operating in a state of transient/no flow (constant pressure), or pulsatile/alternating flow, our work begins to explore the possibilities granted by operating in the closed-loop flow recirculation regime. Here we introduce the concept of FlowBots: soft robots that utilise the characteristics of continuous fluid flow to enable the embodiment of complex control functionality directly into the structure of the robot. FlowBots have robust, integrated, no-moving-part control systems, and these architectures enable: monolithic additive manufacturing methods, rapid prototyping, greater sustainability, and an expansive range of applications. Based on three FlowBot examples: a bidirectional actuator, a gripper, and a quadruped with a swimming gait - we demonstrate how the characteristics of flow recirculation contribute to simplifications in fluidic analogue control architectures. We conclude by outlining our design and rapid prototyping methodology to empower others in the field to explore this new, emerging design field, and design their own FlowBots.
@@ I Introduction
Whilst there as several other notable examples of previous fluidic soft robots that used “pulsatile”, or “cyclic” modes of flow (such as [35, 8, 36, 37, 19, 38]), a condition that has not yet been explored and well-characterised in the context of soft fluidic robots is a state of continuous flow recirculation. Figure 1 illustrates the difference between soft fluidic robots based on the “steady state”, “pulsatile/cyclic”, and “recirculating” modes of flow, as defined in this paper. Fluid can be recirculated in a system for example to achieve thermal regulation, as in the work of Cacucciolo et al. [39], and others [40, 41]. Liao et al. [41] showed that the pressure of this recirculating fluid can also be used to inflate the cavities of an actuator, achieving deformation whilst simultaneously taking advantage of the thermal regulation characteristic (which they used to modify the stiffness of the actuator). An example of a strain gauge based on recirculating fluid can also be found in Koivikko et al. [42], although this came in the form of a separate device, and was not simultaneously used for actuation.
@@
Figure 2: Inspiration for exploiting fluid flow phenomena as a form of embodied intelligence can be found in nature. The geometry of the shark intestine (c) resembles that of a Tesla-valve (a, b). Flow circulation induced by the spiral column geometry results in a diode-like behaviour, making it harder for fluid to flow in one direction (left-to-right as pictured on the figure) than the other (right-to-left). This type of fluidic diode has no moving components; the added functionality is embodied into the structure of the device. Figure (c) has been adapted from the CT scans found in [44].
@@
This paper focuses on embedding additional functionality directly into the bodies of soft robots, by taking advantage of the properties of recirculating fluid flow. Here, we demonstrate that new types of soft robots can be manufactured as a single part, with embedded control, and actuation. This mode of manufacture enables rapid prototyping, increases robustness, reduces waste, and minimises the risk of environmental pollution (opening up opportunities for on-site manufacturing at remote sites, and new applications in e.g. wildlife biodiversity monitoring). We term such robots “FlowBots”.
@@
The focus of this paper is to demonstrate the principles of operation of FlowBots based on three simple examples: a bidirectional actuator, a gripper, and a quadruped robot with a swimming gait. Each example builds upon the operating principles of the previous one. The bidirectional actuator demonstrates how pressure asymmetries arising due to viscous losses can be used for analogue control using localised flow channel constrictions. The gripper shows how flow recirculation can simplify control architectures of multi-actuator systems due to the ability of switching between connection in parallel, and in series. The quadruped FlowBot builds upon the control architecture used for the gripper, and serves as a demonstration of how the embodiment of intelligence in fluid flow characteristics allows to create robots capable of complex behaviours, yet manufactured as single, robust components, without any moving parts. Section II describes some of our experimental work in evaluating the performance of FlowBots with air and water as the operating fluid. We conclude with a discussion on some of the design challenges associated with building FlowBots, empowering others in the field to build their own.
@@ II Results
If the location of the constriction was moved further upstream, or downstream of the tip, the deformation characteristic of the actuator would be permanently skewed to one side. Having the actuator deform differently in both directions at the same pressure difference could be desirable in certain applications. In this paper, all the actuators are symmetric, but we utilise external variable constrictions to regulate the direction of flow, and the pressure difference across the actuator. These constrictions can be implemented using additional valves, by manually blocking/kinking the outlet vents/tubes, or by using tactile, soft buttons (see Figure 6, as well as the Supplementary Video (available at https://wp.me/p9fxLB-Hp). Figure 3 shows how we were able to use two such constrictions to achieve analogue control of the actuator in two directions, by regulating the pressure across it over a continuous spectrum. This analogue mapping can also be achieved by adjusting the direction of flow, as well as the flow rate of the supply pump (hydraulic case), or compressed air supply pressure (pneumatic case).
Figure 3: The principle of operation of the bidirectional actuator. By varying the outlet vent constrictions, the direction, and magnitude of the flow, as well as the pressure difference across the actuator can be controlled with an analogue characteristic. Reversing the flow direction causes the actuator to deform symmetrically in the other direction due to the mirroring of the pressure asymmetry. The percentage value of the constriction is an approximation; we estimated it indirectly by measuring the pressure difference across the actuator using two pressure gauge probes. Note that the flow path from the pressure source to the atmospheric vent is a closed, recirculatory path; this is more intuitive in the hydraulic domain, in which the reservoir, from which the flow/pressure source (PD/centrifugal pump) draws fluid, is only finitely sized.
@@
Section II-A discussed how to control a single bidirectional actuator, two controls (vent outlet constrictions) are used. To fully control two such actuators, four controls would normally be needed. By exploiting the characteristics of recirculating fluid flow, we were able to reduce this fan-in to three. The simplification comes from the fact that under the state of recirculating flow, the actuators can operate not only in a parallel, but also a series connection. The ability to switch between the two configurations allows to simplify the control architecture, and opens up the design space for constructing FlowBots consisting of several actuators, that are capable of more complex behaviours.
@@
The control characteristics of FlowBots are embodied in the interaction of the fluid flow, and the structure of the robot structure. No added parts, or moving components are required, and, as a result, entire robots can be additively manufactured as a single part, in one print job. Figure 5 depicts the operation of a quadruped swimming gait robot that we have built based on the control architecture used for the gripper design described in Section II-B. As in the case of the gripper, we can fully control each pair of limbs for this FlowBot using only three control inputs rather than four, thanks to the ability to switch between parallel and series connection of the actuators. Despite the added complexity of the design, the entire 3D stack of the control architecture comes packaged into the body of the robot, and has been additively manufactured as a single component. After the print job is over, the robot is ready to use. Embodying intelligence in this way makes it easy to prototype FlowBots with ever more complex 3D designs.
Note that the manual connection of the tubing that routs the (same) pressure supply source to the I/O ports on on Figures 3, 4, and 5 is for clarity of illustration only. In Figure 6, we demonstrate how the same task can be accomplished using a tactile, analogue, soft controller that allows a remotely positioned operator to steer the FlowBot by pushing the buttons to adjust the channel restrictions. Please also refer to the Supplementary Video for a depiction of the bidirectional actuator, and swimming gait FlowBots being controlled using the same tactile controller (the former uses a 2-button variant). In an actual application scenario, tethered FlowBots could be controlled in this way, or using electronic valves in place of the buttons.
Figure 5: Operation of a quadruped FlowBot with a swimming gait that builds upon the control architecture used for the gripper described by Section II-B. Due to the embodiment of control into the body of the robot, as well as the characteristics of recirculating flow, FlowBots of increasing complexity can still be manufactured as single components.
Figure 6: Manipulation of an object using the gripper design described by Section II-B. A controller with tactile, soft, buttons, enables the robot operator to control each finger of the gripper independently, by remotely adjusting the flow constrictions using the buttons. This task could also be accomplished via remotely situated (i.e. outside of the operating, potentially extreme, environment) electronic valves.
@@
To account for the difference in the viscosity of the fluids, we used the pressure difference between the inlets of the actuator as the independent variable in the experiment. This approach allowed us to validate whether the actuator deforms the same amount with the same pressure difference, independent of whether air, or water, was used as the working fluid. We examined a range of pressures from 1.25 bar to 2.5 bar (in intervals of 0.25 bar), testing in both flow directions with 3 measurement repetitions per data point. We also repeated the measurements with another identical actuator specimen. Due to the symmetry resulting from placing the channel constriction at the tip of the actuator, we expect it to deform the same amount in both directions, as explained in Section II-A. For more details about the experimental setup, please refer to the Supplementary Material (available at https://wp.me/p9fxLB-Hp).
@@ III Discussion
The channels in the actuators featured in the experiment are as narrow as 0.6 mm (at the tip constriction), and variations on the order of hundreds of microns at different locations along the flow path can disturb the symmetry between the pressure distributions in both directions; such deviations can contribute to marginal differences in the deformation in two directions. In our case, these errors did not significantly affect the robot’s desired functionality, but it might be a factor to consider for those designing FlowBots with very specific precision requirements. In these instances, we recommend carrying out adequate testing to ensure that the desired level of precision has been achieved.
@@
This paper demonstrates how designers can come up with entirely new fluidic soft robot concepts by taking advantage of the characteristics of fluid flow. Constant recirculation of the working fluid comes with the cost of energy expenditure (making the specific systems described in this paper less thermodynamically efficient than their counterparts shown in Figure 1). but it allows to embed additional functionality in the body of the robot, simplify control architectures (as demonstrated in Section II-B), and to rapidly prototype and additively manufacture entirely new FlowBot designs. For additional design considerations to take into account when adapting systems to use recirculating flow, please refer to the Supplementary Material, where we have discussed these points in further detail. For the benefit of others, we have also open-sourced all of the CAD files for the FlowBot, as well as tactile controller, designs that have been showcased in this paper, as well as the Supplementary Video. These files can be accessed on the Soft Systems Group GitHub (https://github.com/orgs/Soft-Systems-Group); the direct link to the repository is also available on the article webpage at https://wp.me/p9fxLB-Hp.
@@
The exact behaviour of the fluid inside of the internal channels is difficult to characterise analytically due to the turbulence arising as a result of the combination of bellow geometries, multi-directional pipe bends, and varying channel cross-sections; this analysis is further complicated by the change in these geometries during actuator deformation, as well as the dynamic interaction of the fluid with the soft material. Whilst we do invite others specialising in the domain to help better characterise these phenomena, we encourage those interested in coming up with their own FlowBot designs to take advantage of Finite Element Method (FEM) software. Despite coming with its own difficulties arising from the same reasons as listed above, it can be a useful tool in helping identify useful trends relating design parameters, and performance. In our design methodology, we have been able to utilise Computational Fluid Dynamics (CFD) to obtain an approximation of the internal flow field that was adequate enough to facilitate effective design exploration and validation (there is also extensive work for others in refining the accuracy of these types of simulations). By combining CFD with structural simulations using Finite Element Analysis (FEA) software we were able to simultaneously explore how modifying parameters in the design affects the fluid flow in the control system, as well as the mechanical deformation of the actuators.
@@
With our chosen method of Fused Deposition Modelling (FDM) 3D printing, we were able to manufacture all of the FlowBots featured in this paper as one part, without any supports or additional post-processing needed. Once taken off the print bed, each FlowBot is ready to use, as it integrates actuation, control, and even fluidic fittings in its design. In spite of the complex bellow geometry of its actuators, and the multi-layered internal 3D control architecture, the swimmer FlowBot described in Section II-C took under 7 hours to print on a cheap, \$200 Artillery Sidewinder X1 printer using Recreus Filaflex 70A filament (as well as 82A for the specimens in Section II-D). This design approach allows for on-site manufacturing at remote sites, is more environmentally sustainable (single material monoliths are easier to dispose of), adds robustness and saves mass and space due to the reduced need for complex routing/manifolds that come with a pressurising/valving system.
@@ IV Conclusion
This paper introduced the concept of FlowBots, a new generation of soft robots that operate under a state of fluid flow recirculation. FlowBots are controlled by taking advantage of fluid flow characteristics, allowing for embodied intelligence. This work demonstrates the operation of three simple FlowBot systems: a bidirectional actuator, a gripper, and a quadruped robot with a swimming gait.